If the measure of two exterior angles, one at each vertex of a triangle is 230, find the measure of the third exterior angle and its adjacent interior angle
Answer by addingup(3677)

The sum of the exterior angles for any figure no matter how many sides is equal to 360. In your problem:
360-230 = 130 is the 3rd exterior angle.
`````````````````````````````````````````````````````
Exterior Angle Theorem
The measure of an exterior angle of a triangle equals the sum of the measures of the two remote interior angles of the triangle. See my illustration below, W = X+Y.
So, in your problem X+Y= 130
therefore, angle Z = 180-130 = 60